### v1

# Should you use LLMs to simulate opinions? Quality checks for early-stage deliberation

Abstract
The array of emergent capabilities of large language models (LLMs) has sparked interest in assessing their ability to simulate human opinions in a variety of contexts, potentially serving as surrogates for human subjects in opinion surveys. However, previous evaluations of this capability have depended heavily on costly, domain-specific human survey data, and mixed empirical results about LLM effectiveness create uncertainty for managers about whether investing in this technology is justified in early-stage research. To address these challenges, we introduce a series of quality checks to support early-stage deliberation about the viability of using LLMs for simulating human opinions. These checks emphasize logical constraints, model stability, and alignment with stakeholder expectations of model outputs, thereby reducing dependence on human-generated data in the initial stages of evaluation. We demonstrate the usefulness of the proposed quality control tests in the context of AI-assisted content moderation, an application that both advocates and critics of LLMs’ capabilities to simulate human opinion see as a desirable potential use case. None of the tested models passed all quality control checks, revealing several failure modes. We conclude by discussing implications of these failure modes and recommend how organizations can utilize our proposed tests for prompt engineering and in their risk management practices when considering the use of LLMs for opinion simulation. We make our crowdsourced dataset of claims with human and LLM annotations (TopicMisinfo) publicly available for future research.
1 Introduction
Large language models (LLMs) have demonstrated significant appeal and versatility across a wide range of tasks. While the primary training objective of the base model is to predict the next most likely token, training on vast and diverse datasets, combined with increasingly complex system architectures has resulted in the “emergence” of capabilities that were not explicitly anticipated during development [1]. For instance, LLMs have been shown to learn new tasks through in-context learning [2], perform coding tasks with remarkable accuracy [3], and have displayed human-like capabilities across different tasks [4, 5].
These emerging capabilities have led researchers to investigate more unconventional uses of LLMs, including automating opinion surveys by simulating the viewpoints of specific demographic or ideological groups [6]. The possibility that LLMs may successfully simulate demographic viewpoints stems from (i) evidence that LLMs can be prompted to replicate responses from human subject studies in psychology [7], (ii) training data that likely encompass a broad spectrum of perspectives and opinions [8], and (iii) prompting techniques that lead LLMs to produce seemingly coherent, opinionated responses [9, 10]. Successfully automating such surveys could have far-reaching implications in both private and public sectors, ranging from marketing [11] and content moderation [12] to policymaking and public relations [13]. Researchers have noted that LLMs for opinion simulation could be particularly valuable for pilot studies [14, 15, 16], for sampling opinions of notoriously hard-to-reach populations [17], and for annotating data that can be psychologically harmful for human subjects [18].
However, existing empirical findings are mixed regarding the suitability of LLMs for human opinion simulation. Some studies indicate that LLMs prompted with specific demographic characteristics produce accurate predictions of group-level opinions [16, 19], while other studies show that LLMs prompted with demographic characteristics can fail to adequately represent group-level opinions, reporting particularly on inadequate performance for minority groups and non-Western identities [9, 20, 6, 21]. Findings have also shown that LLMs tend to flatten within-group differences of opinion [18, 22, 21], and may not generalize across different opinion tasks and domains [13, 23].
These mixed results indicate that, while there is a possibility that current or future iterations of LLMs could be leveraged to successfully simulate opinions—a particularly valuable prospect when surveys expose human subjects to harmful content [18] (e.g., content moderation) or trade secrets (e.g., marketing new products)—thorough evaluation is needed before relying on LLM outputs for any such application. However, current methodologies to evaluate LLM-based opinion simulation rely heavily on costly human data [16, 19, 9, 20, 18]. These costs present challenges to initial research and development efforts of LLMs for automating opinion surveys for two main reasons. First, given the existing uncertainty around whether LLMs effectively simulate opinions, there may be little incentive to commit financial resources up front. Second, paradoxically, if using simulated opinions in a pilot study requires human data to validate any conclusions, one may as well simply use human data. These challenges are exacerbated by the dynamic evolution of human opinion, as well as LLM-generated ones due to system updates.
To address the challenge of costly reliance of existing validation methodologies on human data, we develop a series of tests for early-stage deliberation that do not require human data for evaluation and instead probe LLM outputs for other desirable qualities. Specifically, we propose tests to evaluate whether LLMs demonstrate:
  1. 1.
    Logical Consistency—whether models position “average” or “neutral” opinions as a convex combination of group-level opinions;
  2. 2.
    Model Stability Across Updates—whether models remain relatively consistent in their opinion outputs over time across minor updates;
  3. 3.
    Alignment with Stakeholder Expectations—whether models position differences in group opinions in ways that are consistent with prior domain knowledge.
These tests are informed by research on assessing quality standards and biases in subjective data annotation [24, 25, 26, 27, 28, 29, 30, 31], and they are intended to help practitioners decide whether additional investment in human data is justified and guide the iterative design of prompts intended to meet these quality standards. We show that these proposed quality control tests, despite their simplicity, reveal crucial failure modes of state-of-the-art, commercially available LLMs, indicating that these are not yet reliable means to simulate group-level opinions. We conclude by discussing implications of the observed failure modes and recommend how organizations can utilize our proposed tests for prompt engineering and in their risk management practices when considering the use of LLMs for opinion simulation. We make our crowdsourced dataset of claims with human and LLM annotations (TopicMisifo) publicly available for future research.
Experimental Dataset
Content moderation is often cited as an appealing application for LLMs. This is because moderation tasks can be psychologically harmful for human workers [32], occur at high frequency [33], and demand the scalability that automated systems can provide [34]. We explore automating opinion surveys in the context of fact-checking workflows, where LLM-based opinion surveys could help scale how claims are prioritized for human verification. Similar to many other survey contexts, in this content moderation tasks the social background and demographic identities of human annotators can crucially shape their assessments [30, 35, 36]. It is critical that AI-driven systems should capture relevant variations in opinions across demographic groups, which can reflect a range of group-specific priorities and perspectives. Otherwise, it risks systematically overlooking the concerns of certain demographic groups, leaving them disproportionately vulnerable to misinformation harms [37]. Indeed, the risk of overlooking harms to certain demographic groups has already been observed among fact-checkers employing advanced claim prioritization tools and technologies [38, 39, 40].
To provide an empirical application of the proposed evaluation methodologies, we created the TopicMisinfo dataset, which consists of 160 claims—some true, some false—spanning a wide range of topics. These include potentially demographically divisive issues such as abortion and illegal immigration, as well as topics like entertainment and weather, which are generally associated with broader consensus. Anchored on results from a survey from the American National Election Studies (ANES) [41], we selected topics based on their likelihood to provoke disagreement or agreement about whether they warrant fact-checking, considering four key criteria: (i) perceived importance for prioritization, (ii) general public interest, (iii) potential harm to specific groups, and (iv) perceived factual validity. The focus of our study is on evaluating LLMs ability to simulate gender perspectives on claims on the third axis — potential harm of claims to specific groups — as research and past polling has pointed out that gender differences tend to be particularly salient for this item [42, 43].
We apply the proposed evaluation framework to a diverse set of commercially available LLMs: GPT-3.5-Turbo (checkpoints 06-13-23, 11-06-23, and 01-25-24), GPT-4, LLaMA 3, Titan-Text-Premier, and Mistral-Large—hereafter referred to as GPT3.5a, GPT3.5b, GPT3.5c, GPT-4, LLaMA3, Titan, and Mistral, respectively. To simulate group-level responses, we adopt the “backstory prompting” methodology [16], instructing models to respond from the perspectives of either men or women. These prompts build on established techniques for gender-specific language model conditioning [16, 20]. Full prompt details are available in Supplementary Materials S3. To assess alignment with stakeholder expectations—one of our three proposed evaluation criteria—we use two complementary strategies. First, we incorporate “gold standard” items inspired by comprehension-based quality control protocols in human annotation [27] to evaluate baseline output reliability for this task in Experiment 3a. Second, we incorporate crowdsourced annotations to demonstrate how human data—when available—can inform domain knowledge and serve as a valuable comparison point, which we address in Experiment 3b. See Supplementary Materials S1 and S2 for further information on dataset construction, examples of claims used in the opinion survey, and descriptive statistics of human annotations.
2 Results
In this section, we analyze the results of three quality checks for a particular content moderation task: predicting the harm of potential misinformation to certain groups as a means to prioritize it for review [^1] While not the only criteria relevant for prioritization, research has noted that this is an important axis of consideration used by fact-checkers [44].. These tests, designed to emphasize logical consistency, model stability, and alignment with stakeholder expectations, serve as sanity-checks in early-stage deliberation when considering the use of LLMs to simulate group-level opinions.
Experiment 1: Logical Consistency
Figure 1: Experiment 1 Results. Plausible ranges of $q_{0}$ across topics for different models are highlighted in blue.
Testing for logical consistency can be an effective way to gauge the reliability of human annotators [29] or data aggregation methods [45]. For instance, Arguete et al. [29] use logical contradiction as a measure of human annotator inattention. We introduce a related methodology for evaluating logical consistency of LLMs in group-level opinion simulation by testing whether a model, when prompted for an “average" or “neutral” perspective on a Likert-scale assessment, generates a stance that is a convex combination of its predicted group-level opinions. If a model’s average opinion response is positioned plausibly within the convex hull of divergent group opinions, it suggests a structured and internally consistent approach to generating viewpoints.
In this paper, we define statistical neutrality as an opinion located within the convex hull of all group-level opinions within a social category (e.g., gender, race, etc.), including the extrema (i.e., only representing one group). This is distinct from neutrality focused on disinterest or impartiality [46, 47]. However, there is no single established approach to prompting an LLM to generate a “statistically neutral” viewpoint. To address this, four distinct prompts were evaluated, each with different approaches to querying statistical neutrality. See Supplementary Materials S3 for exact prompt specifications.
To determine whether these neutral prompts position responses between men’s and women’s predicted opinions, we introduced a mixture parameter $\hat{q}$ that interpolates between the model’s predicted opinion for women ($\hat{q}=0$) and men ($\hat{q}=1$) on a given claim. If an LLM’s neutral response yields $\hat{q}\in[0,1]$, it suggests that the model is, at least to some degree, operating in a logically consistent manner. Further methodological details are provided in the Materials and Methods section.
The results, presented in Figure 1, reveal several patterns across models and topics.
GPT3.5c
Base Prompts 1, 2, 3, and 4 were effective in generating statistically neutral opinions on the divisive topic of abortion, with feasible $q_{0}\in[0,0.85]$ for Base Prompts 3 and 4 and $q_{0}\in[0,0.55]$ for Base Prompts 1 and 2. No feasible $q_{0}$ was found for the topics Black Americans, Illegal Immigration, and LGBTQ issues, indicating inconsistent positioning of statistically neutral perspectives for these topics.
LLaMa3
All prompts successfully produced statistically neutral opinions for the topic of Black Americans (plausible $q_{0}\in[0,0.7]$), and Base Prompts 1 and 4 additionally generated statistically neutral (though men-leaning) opinons for illegal immigration (plausible $q_{0}\in[0.85,1]$). No plausible $q_{0}$ was found for abortion or LGBTQ issues across prompts.
Mistral
This model failed to demonstrate feasible $q_{0}$ values across all divisive topics and prompts, suggesting significant limitations in its ability to provide internally consistent perspectives.
Titan
Base Prompt 4 alone generated plausible $q_{0}\in[0,1]$ for the topic Black Americans. Base Prompts 1 and 2 demonstrated a broad range of plausible $q_{0}$ values for the LGBTQ topic ($q_{0}\in[0,1]$) but Base Prompts 3 and 4 did not produce a feasible $q_{0}$, highlighting significant prompt sensitivity. No feasible $q_{0}$ was found for topics Abortion and Illegal Immigration.
Experiment 2: Model Stability Over Time
Measuring the stability of a single LLM’s outputs across minor model updates (or “checkpoints”) offers critical insight into its reliability. This kind of test is motivated by quality control in human annotation pipelines, where inconsistent judgments cast doubt on an annotator’s credibility [28]. As Abercrombie et al. [25, pg.1] observe, “consistency and label stability are important because, without them, annotation schemes are unlikely to be repeatable or reproducible.”
Model stability is also closely tied to the concept of backwards compatibility, a long-standing concern in software engineering [48], and one of growing importance when we consider the use of “off-the-shelf” machine learning models applied to a variety of application domains [49]. In AI systems, backwards compatibility refers to whether updated models make categorically similar predictions or errors to previous versions. For LLMs, it is reasonable to expect backwards compatibility within a given “model class” (e.g., across checkpoints of GPT-3.5-Turbo), though not necessarily between model classes (e.g., GPT-3.5-Turbo vs. GPT-4). Updates within a model class are generally intended to improve safety and reliability, not to introduce significant behavioral or quality changes [50]. Yet, emerging evidence suggests this expectation is not always met. For example, Chen et al. [51] found that two checkpoints of GPT-3.5-Turbo disagreed on 27% of multiple-choice responses in the OpinionQA dataset.
In this experiment, we introduce a statistical test to assess whether a model class maintains relatively stable group-level opinion outputs over time. Specifically, we evaluate how Likert scale responses vary across checkpoints within the same model class for men-conditioned, women-conditioned, and gender-neutral prompts. We report our test statistic ($\hat{L}$) and corresponding p-values as defined in the Materials and Methods section. Figure 2 visualizes the mean rating changes across three GPT-3.5-Turbo checkpoints, broken down by topic and prompt condition. Notably, we observe that both quality control items (e.g., Gold items from the TopicMisinfo dataset) and divisive topics exhibit statistically significant fluctuations, highlighting key areas of instability that may affect the reliability of LLM-generated opinions over time.
Figure 2: Experiment 2 Results. This chart highlights statistically significant changes in model outputs across three updates (June ‘23, Nov ‘23, and Jan ‘24) for three types of prompts: neutral, men-conditioned, and women-conditioned.
Gold
Ratings of these quality control items should remain stable and low over time. These claims were all designed to be harmless truths or lies. However, we see significant differences in ratings across checkpoints for men-conditioned ($\hat{L}_{\text{men}}$(GPT3.5a, GPT3.5c) = 0.19, p $<$ 0.001), neutral ($\hat{L}_{\text{neutral}}$(GPT3.5a, GPT3.5c) = 0.55, p $<$ 0.001), and women-conditioned prompts ($\hat{L}_{\text{men}}$(GPT3.5a, GPT3.5c) = 0.51, p $<$ 0.001).
Abortion
Men-conditioned and neutral prompts showed substantial shifts in ratings from the first to the second checkpoint ($\hat{L}_{\text{men}}$(GPT3.5a, GPT3.5b) = 1.25, p $<$ 0.001; $\hat{L}_{\text{neutral}}$(GPT3.5a, GPT3.5b) = 0.81, p $<$ 0.001), while women-conditioned prompts remained stable ($\hat{L}_{\text{women}}$(GPT3.5a, GPT3.5b) = 0.11, p = 0.46). There was no significant change from the second to third checkpoint for any gender condition.
Claims about Black Americans
For the neutral and men-conditioned prompts, we see significant change from the first to the second checkpoints ($\hat{L}_{\text{neutral}}$(GPT3.5a, GPT3.5b) = 1.54, p $<$ 0.001; $\hat{L}_{\text{men}}$(GPT3.5a, GPT3.5b) = 2.05, p $<$ 0.001). Ratings increased significantly across all prompt conditions from the first to the final checkpoint ($\hat{L}_{\text{men}}$(GPT3.5a, GPT3.5c) = 2.05, p $<$ 0.001; $\hat{L}_{\text{women}}$(GPT3.5a, GPT3.5c) = 0.50, p $<$ 0.001; $\hat{L}_{\text{neutral}}$(GPT3.5a, GPT3.5c) = 1.21, p $<$ 0.001).
Illegal Immigration
Neutral and men-conditioned prompts showed significant increases in ratings over time ($\hat{L}_{\text{neutral}}$(GPT3.5a, GPT3.5c) = 1.82, p $<$ 0.001; $\hat{L}_{\text{men}}$(GPT3.5a, GPT3.5c) = 0.39 , p $<$ 0.001), while women-conditioned prompts showed a statistically significant decline ($\hat{L}_{\text{women}}$(GPT3.5a, GPT3.5c) = 1.05, p $<$ 0.001).
LGBTQ Issues
Responses varied markedly by prompt condition. Neutral prompts exhibited increased sensitivity to claims across updates ($\hat{L}_{\text{neutral}}$(GPT3.5a, GPT3.5b) = 0.56, p $<$ 0.001; $\hat{L}_{\text{neutral}}$(GPT3.5b, GPT3.5c) = 0.92, p $<$ 0.001). Men-conditioned prompts showed a large increase followed by a decrease ($\hat{L}_{\text{men}}$(GPT3.5a, GPT3.5b) = 1.86, p $<$ 0.001; $\hat{L}_{\text{men}}$(GPT3.5b, GPT3.5c) = 1.07, p $<$ 0.001). Women-conditioned prompts remained stable throughout.
Experiment 3: Alignment with Stakeholder Expectations
Table 1: Salient Examples of Opinion Differences Diverging from Priors in Experiment 3a. $\mu(c_{i})^{gender}_{LLM}$ represents the average predicted harmfulness rating for the claim for a given gender.
Claim Prior Model $\mu(c_{i})^{men}_{LLM}$ $\mu(c_{i})^{women}_{LLM}$ Difference A circle is round. Not divisive GPT3.5c 1 1.71 -0.71 All diseases can be spread through radio waves Not divisive Titan 1 6 -5 An almond is a type of nut Not divisive GPT3.5c 1 2.25 -1.25 Diamonds are very expensive Not divisive GPT3.5c 1 3.14 -2.14 You’ll never be able to have children if you have an abortion Divisive LLaMa 6 6 0 The Department of Defense pays for troops’ abortions Divisive Titan 1 1 0 Minors can receive abortions without parental consent in Minnesota Divisive Mistral 5 5 0 Georgia Politician Stacey Abrams Suggested Abortions Can Ease Inflation Divisive LLaMa 4 4 0
Table 2: Experiment 3b Statistical Results. Significant positive results reflect a tendency to exaggerate differences between men and women, while significant negative results reflect a tendency to erode differences where they actually exist.
Topic gpt-35-23-0613 gpt-35-23-1106 gpt-35-24-0125 gpt-4 llama3 mistral-large titan-text-premier
$\hat{L}_{\omega}$ p-val $\hat{L}_{\omega}$ p-val $\hat{L}_{\omega}$ p-val $\hat{L}_{\omega}$ p-val $\hat{L}_{\omega}$ p-val $\hat{L}_{\omega}$ p-val $\hat{L}_{\omega}$ p-val
More Divisive Topics
Abortion 1.14 *** 0.10 0.32 -0.89 *** -0.89 *** -0.69 *** -0.65 ***
Black Americans 0.88 *** -0.08 -0.14 -0.84 *** -0.65 *** -1.15 *** -0.59 **
Illegal Immigration 0.56 *** -0.56 ** -0.37 -0.75 *** -0.44 * -0.37 -0.40 *
LGBTQ 0.51 ** -0.72 *** -0.15 -0.44 * -0.53 ** -0.49 * -0.53 **
Less Divisive Topics
Entertainment 0.84 *** 0.49 ** 0.56 *** 0.18 0.20 0.28 *** 0.25 *
Gold 0.25 *** 0.19 *** 0.37 *** -0.00 -0.14 0.01 0.19 ***
HealthScience 1.05 *** 0.52 * 0.38 -0.10 0.02 0.38 * 0.31 *
Sports 0.01 -0.01 0.13 -0.11 -0.11 -0.11 -0.06
USA 1.09 *** 0.58 *** 0.36 ** -0.31 ** -0.09 -0.22 * 0.14
WeatherClimate 0.65 ** 0.62 ** 0.46 * -0.18 0.02 -0.34 ** 0.07
Significance levels: * for p $<$0.05, **∗ ∗ for p $<$ 0.01, ***∗ ∗ ∗ for p $<$ 0.001.
Blue: Exaggeration compared to human data, Orange: Erosion compared to human data.
Stakeholders that rely on LLM-simulated opinions for decision making have a vested interest in verifying that LLMs accurately reflect group-level viewpoints. If an LLM does not position group-level opinions appropriately, it risks introducing biases into human decision-making that depends on these model outputs. For example, if fact-checking systems rely on LLMs to estimate which claims different groups are likely to find harmful, a model that underrepresents gender differences in concern may deprioritize fact-checks on topics that disproportionately affect one group’s information environment. Notably, even in the absence of “ground truth" human data to validate a particular opinion survey task, stakeholders often have valuable domain knowledge that can be leveraged to validate the reliability of LLMs. In this experiment, we focus on the positioning of gender differences in opinion, asking: Do models position men’s and women’s opinions in ways that align with stakeholder expectations regarding the degree of disagreement across different topics?
We define “expectations” by leveraging topics where prior research, survey results, or common sense strongly indicates minimal or substantial disagreement between gender groups. Experiment 3 examines two complementary scenarios: (a) whether the model violates well-established priors from common sense, domain knowledge, or survey data—e.g., that a neutral statement like “A circle is round” should not elicit divergent opinions by gender, whereas a claim about abortion might, drawing inspiration from prior comprehension-based quality control checks [27]; and (b) whether the model distorts stakeholder expectations on less obviously gender-polarized issues (e.g., politics, immigration, health and science), by introducing task-specific human survey data that serves as a reasonable prior for stakeholder expectations. This two-part design allows us to identify when a model introduces artificial rifts or masks genuine ones, both crucial factors in gauging an LLM’s trustworthiness for representing group-specific viewpoints. See Materials and Methods for details on the methodology for Experiments 3a and 3b.
Exaggeration of Opinion Differences on Expected Consensus Topics
Experiment 3a
To test for exaggeration of opinion differences, we compared the output of LLMs prompted to reflect the opinions of men and women on topics where we had very strong priors that no division should exist. For instance, the Gold topic in the TopicMisinfo dataset contains a set of items that are harmless and uncontroversially true (e.g., “A circle is round”) or false (e.g., “The tallest tree on Earth creates a land bridge to Mars”), and they were compiled and used as a means to make sure that human annotators were paying attention and comprehending the task. We would very reasonably expect no opinion disagreement on claims of this type. However, we see GPT3.5a ($\hat{L}_{GPT3.5a}=0.25$, p $<$ 0.001), GPT3.5b ($\hat{L}_{GPT3.5b}$ = 0.20, p $<$ 0.05), and GPT3.5c ($\hat{L}_{GPT3.5c}$ = 0.38, p $<$ 0.01), as well as Titan ($\hat{L}_{Titan}$ = 0.19, p $<$ 0.01), tended to produce statistically significant differences in opinions by gender for claims on this topic. See Table 1 for specific examples of exaggeration.
Experiment 3b
When preliminary data from real human responses is available, this can inform the stakeholder prior. For this purpose, we use crowdsourced annotations of the TopicMisinfo dataset. When comparing LLM responses to crowdsourced human responses, GPT3.5a, GPT3.5b, and GPT3.5c tended to exaggerate differences between groups, particularly for topics that are not typically considered divisive between genders. For example, claims from topics such as entertainment ($\hat{L}_{GPT3.5a}=0.84$, $p<0.001$; $\hat{L}_{GPT3.5b}=0.49$, $p<0.01$; $\hat{L}_{GPT3.5c}=0.55$, $p<0.001$), health & science ($\hat{L}_{GPT3.5a}=1.05$, $p<0.001$; $\hat{L}_{GPT3.5b}=0.52$, $p<0.01$), U.S. politics ($\hat{L}_{GPT3.5a}=1.09$, $p<0.001$; $\hat{L}_{GPT3.5b}=0.58$, $p<0.001$; $\hat{L}_{GPT3.5c}=0.36$, $p<0.01$), and weather & climate ($\hat{L}_{GPT3.5a}=0.65$, $p<0.01$; $\hat{L}_{GPT3.5b}=0.62$, $p<0.01$; $\hat{L}_{GPT3.5c}=0.46$, $p<0.05$) saw significant overestimation of group-level disagreement by these models.
Erosion of Opinion Differences on Expected Divisive Topics
Experiment 3a
To test for erosion of opinion differences, we compared the output of LLMs prompted to reflect the opinions of men and women on topics where we had very strong priors that division should exist. For instance, it is well-established that men and women may perceive harms differently by claims about abortion and women’s health [43, 42], and this topic is included in the TopicMisinfo dataset. Thus, we have some prior that a reasonable LLM should display differences in outputs when conditioned to respond as if it were either a man or woman. However, we didn’t find any significant differences between the outputs from men- or women-conditioned prompts on the topic for Mistral ($\hat{L}_{Mistral}=0.26$, p = 0.09), GPT-4 ($\hat{L}_{GPT4}=0.06$, p = 0.16), and LLaMa3 ($\hat{L}_{LLaMa}=0.06$, p = 0.23). See Table 1 for specific examples of erosion.
Experiment 3b
We saw that GPT-4, LLaMa3, Mistral, and Titan tended to erode differences, underestimating the degree of disagreement in topics where human data suggest significant divergence. For instance, these models consistently underestimated the extent to which women expressed more concern than men about abortion-related claims ($\hat{L}_{GPT4}=-0.89$, $p<0.001$; $\hat{L}_{LLaMa}=-0.89$, $p<0.001$; $\hat{L}_{Mistral}=-0.69$, $p<0.001$; $\hat{L}_{Titan}=-0.65$, $p<0.001$), claims about Black Americans ($\hat{L}_{GPT4}=-0.84$, $p<0.001$; $\hat{L}_{LLaMa}=-0.65$, $p<0.001$; $\hat{L}_{Titan}=-0.65$, $p<0.001$), and claims about LGBTQ related issues ($\hat{L}_{GPT4}=-0.44$, $p<0.05$; $\hat{L}_{LLaMa}=-0.53$, $p<0.01$; $\hat{L}_{Mistral}=-0.49$, $p<0.05$; $\hat{L}_{Titan}=-0.53$, $p<0.01$).
3 Discussion
Quality control methods assist in early-stage deliberation
LLMs offer a compelling approach to simulating demographic opinions, particularly because they can generate plausible outputs without requiring any external data to get started. However, assessing whether these outputs faithfully capture the nuanced patterns of human perspectives across diverse groups, tasks, and topics remains a significant challenge. This difficulty creates space for premature or unsupported claims about their reliability. The problem is compounded by the fact that LLMs rarely refuse to answer; instead, they often produce responses that appear coherent and reasonable on the surface, even when the underlying logic or representational accuracy is flawed. In this paper, we propose a set of conceptual checks—focused on logical consistency, model stability, and alignment with basic stakeholder expectations—to provide a principled initial foundation for evaluating whether LLM-simulated opinions merit further investment and can inform prompt engineering efforts early in the research and development timeline.
Applying these methods to several state-of-the-art models, we provide empirical evidence that these models fail to yield reliable results across our three checks. These observed shortcomings can mislead decision-makers by distorting genuine opinion distributions and skewing outcomes in contexts such as market research, policy development, and strategic planning. Our findings underscore the need for caution and careful prompt engineering when deploying these technologies.
The proposed methodology serves as a practical framework for early-stage deliberation, equipping organizations with tools to assess LLM reliability before investing in large-scale human reference data. By integrating these quality checks, organizations can make more informed decisions about the deployment of LLMs and implement adequate risk management strategies. In the following sections, we analyze observed failure modes from our three experiments, their consequences, methodological implications, and provide insight into how organizations can use these tests to inform and manage the associated risk of these technologies.
Implications: Logical Inconsistency of Neutral Opinions
Observed Failure Modes and Consequences
Experiment 1 reveals that most models fail to produce logically consistent neutral responses: although a neutral stance should logically be represented as a convex combination of the model’s predicted opinions for different demographic groups, only 20% of the 64 model–prompt–topic combinations passed this check. The remaining 80% produced “neutral” outputs that were actually more extreme than both men’s and women’s predicted opinions, casting doubt on whether these models reflect genuine opinion distributions or merely artifacts of their training. Such inconsistencies undermine the idea that opinion estimation might be a legitimate “emergent” LLM capability, thereby raising concerns about these models’ suitability for opinion surveys.
Methodological Implications
Logical consistency is a cornerstone of annotation quality: outputs that contradict each other without justification reduce trust in the annotator, whether human or machine. When LLMs violate basic structural expectations—such as the expectation highlighted in this experiment—it suggests a breakdown in internal reasoning. This failure challenges the use of LLMs as annotation tools or survey proxies. Our findings point to a critical gap: LLM evaluations need to account for internal relational consistency between related prompts, not just isolated outputs. Without this kind of internal logic check, high-quality-sounding responses may mask logical errors.
Risk Management
Although failing this check raises broader doubts about an LLM’s capacity to model group-level opinions, its practical impact depends on how central a coherent notion of neutrality is to the task at hand. If a use case primarily relies on modeling group-specific opinions, flawed neutral outputs may not be critical. However, for tasks like gauging overall sentiment or developing balanced recommendations, this failure mode becomes more pertinent. When these issues arise, multi-stage prompts that incorporate previously predicted group-level opinions as context might remedy them, allowing LLMs to position neutrality in their own context.
Implications: Unstable Outputs Across Model Checkpoints
Observed Failure Modes and Consequences
Experiment 2 shows that even successive versions of the same model can exhibit unstable outputs, implying a lack of backwards compatibility of prompts. This instability complicates tasks such as policymaking, content moderation, and market research, where stable, interpretable opinions are critical. Third-party LLM APIs exacerbate the problem: providers frequently modify models to fix biases or improve performance, yet these opaque updates can invalidate prior workflows and expectations [50, 52].
Methodological Implications
Our findings highlight the importance of evaluating not just a single model, but also comparing different checkpoints (and, potentially, related models) to assess reliability and backwards compatibility. When outputs shift significantly between model versions, and there is no consistent sampling of human opinions over time, it becomes difficult to determine whether these changes reflect genuine shifts in public attitudes or are instead driven by exogenous factors, such as new training data, updated reinforcement learning objectives, or other changes in the model pipeline. This uncertainty poses a particular challenge for applications that rely on stability of decision-making criteria over time, such as content moderation. Moreover, unlike traditional software, LLMs often lack transparent, version-locked updates, reinforcing the need for continuous evaluation rather than one-time benchmarking.
Risk Management
Dedication to using a version-controlled model can help reduce some of the instability. Regardless, selecting the “best” model and prompt under these conditions is challenging: newer models are not guaranteed to outperform older ones since this task is certainly not directly optimized for, and improvements on one front (e.g., harmlessness or safety) may significantly impact prompt performance. While ensembling models can, in principle, mitigate uncertainty [53], shared training data and alignment constraints between models from the same model class may leave their biases too closely correlated to observe the usual benefits. Practitioners should test is whether in-context response calibration can offset the variability noted in our experiments. Future work should investigate which forms of contextual scaffolding—such as embedding representative edge cases or eliciting chain-of-thought reasoning—most effectively stabilize model outputs across checkpoints.
Implications: Failure to Meet Stakeholder Expectations
Observed Failure Modes and Consequences
Experiment 3 reveals systematic failures in how LLMs position group-level opinions. Older models like GPT-3.5-turbo often exaggerate opinion differences, even on common-sense, non-divisive claims, such as “An almond is a type of nut” and “A circle is round,” risking stereotyping and misallocation of resources in downstream applications like claim prioritization. Newer models (GPT-4, LLaMa3, Mistral, Titan), by contrast, frequently erode meaningful differences, particularly on topics known to be gender-divisive. This compression of diversity can obscure key priorities for certain groups. This may reflect an implicit refusal where models respond superficially but sidestep the task of representing nuanced disagreement.
We also observed inconsistent error direction across models. For the same topic (e.g., immigration), one model (e.g., GPT3.5a) may exaggerate differences while another (e.g. GPT-4) erodes them. These contradictions, combined with uncertainty in the human reference data itself, introduce irreducible uncertainty in model selection: we cannot definitively say which model is “closer to the truth” without assuming the reference data is unbiased or that one model is inherently more reliable than the other, both of which are challenging claims to validate in the context of opinion simulation.
Methodological Implications
Our approach—using curated topics in the TopicMisinfo dataset with strong societal priors—provides a domain-grounded alternative to noisy or expensive human reference data. If a model positions large gender gaps on statements like “A circle is round,” or erases known gender divides on issues like abortion, it likely fails to capture basic structures in group sentiment, and all of its outputs should be considered suspect. These tests serve a function analogous to inter-annotator agreement in human annotation pipelines [54, 26, 55]. Rather than treating human data as absolute ground truth, we assess whether models reflect well-established social patterns. This allows for lightweight, reproducible tests of model reliability without large-scale annotation efforts.
Risk Management
Organizations should prioritize testing whether model outputs align with clear expectations in their domain. Errors on divisive topics may reflect real-world complexity, but fabricating disagreement on neutral issues specifically designed for quality control can yield a high-quality signal that model performance is unreliable. In settings without established priors, domain experts can still construct targeted edge cases to test model reasoning. Including system instructions or information in the prompt that draws attention to the risks of both exaggeration and erosion of group differences may help calibrate responses with stakeholder expectations. Additionally, chain-of-thought prompts can be used to surface uncertainty or re-position group opinion estimation after an initial prediction [56].
Conclusion
By applying quality control checks that surface logical consistency, LLM self-consistency over time, and stakeholder expectations of model performance, practitioners can determine whether an LLM clears a basic threshold for reliability before devoting significant resources to collecting large-scale human reference data to fine-tune model performance. A model that meets these conditions may still require subsequent testing and tuning, but it at least demonstrates fundamental viability. If it fails on multiple fronts, stakeholders should carefully consider whether an LLM-driven approach is worthwhile or whether alternative data-centric or human-led solutions would be more consistent and trustworthy.
4 Materials and Methods
General Bootstrap Hypothesis Testing Framework
All experiments in this study rely on a common bootstrap hypothesis-testing framework, following well-established guidelines [57, 58]. Although the test statistic and specific sampling procedures differ by experiment, the overall structure is as follows:
  1. 1.
    Define the test statistic. We specify a test statistic $L$ that measures the relevant difference or relationship of interest (e.g., the divergence between LLM and human responses, or between outputs from two LLM checkpoints).
  2. 2.
    Construct a pooled set under the null hypothesis. The null hypothesis $H_{0}$ varies by experiment but generally posits “no difference” in the phenomenon under investigation. Under $H_{0}$, we pool all annotations (or responses) across the conditions being compared, reflecting the assumption of no effect.
  3. 3.
    Bootstrap resampling. We generate $B=10{,}000$ bootstrap samples by resampling with replacement from the pooled set, preserving each condition’s original sample size. Each bootstrap sample yields a recalculated value of the test statistic, $L^{b}$.
  4. 4.
    Compute the p-value. Because all of our experiments use two-sided tests, the p-value is defined as the proportion of bootstrap replicates for which the absolute value of the bootstrapped statistic exceeds that of the observed statistic. Formally:
    $T=\frac{1}{B}\sum_{b=1}^{B}I\bigl{[}\bigl{|}\hat{L}^{b}\bigr{|}>\bigl{|}\hat{L}\bigr{|}\bigr{]},$ (1)
    where $\hat{L}$ is the observed (empirical) test statistic, and $\hat{L}^{b}$ is the statistic in the $b$-th bootstrap sample. We reject $H_{0}$ if $T<\alpha$, where $\alpha$ is the chosen significance level.
In the sections below, we describe how this framework is adapted to each experiment’s unique test statistic, resampling approach, and research questions.
Experiment 1 Methodology
This experiment examines whether large language models plausibly position neutral responses as convex combinations of group-level responses. In other words, we test whether a neutral response can be expressed as
$\mu\bigl{(}c_{i}\bigr{)}_{\text{LLM}}^{\text{neutral}}\;=\;\sum_{g=1}^{G}q_{g}\,\mu\bigl{(}c_{i}\bigr{)}_{\text{LLM}}^{g}\quad\begin{array}[]{l}\text{subject to:}\\ \sum_{g=1}^{G}q_{g}=1,\\ q_{g}\geq 0\;\;\text{for all }g,\end{array}$ (2)
where $G$ is the number of demographic groups within a social category, and $\mu\bigl{(}c_{i}\bigr{)}$ is the mean of the empirical distribution of responses. We consider two groups (men and women) within the social category of gender (acknowledging the limitation of not including non-binary genders). Equation 2 thus simplifies to:
$\mu(c_{i})_{LLM}^{\text{neutral}}=q\times\mu(c_{i})_{LLM}^{\text{men}}+(1-q)\times\mu(c_{i})_{LLM}^{\text{women}},\quad q\in[0,1].$
To determine how the LLM’s actual neutral responses compare to this idealized mixture, we estimate the implied mixture weight for each claim $c_{i}$:
$\hat{q}_{c_{i}}=\frac{\mu(c_{i})_{LLM}^{\text{neutral}}-\mu(c_{i})_{LLM}^{\text{women}}}{\mu(c_{i})_{LLM}^{\text{men}}-\mu(c_{i})_{LLM}^{\text{women}}},$ (3)
where $\mu_{m}=\mu(c_{i})_{LLM}^{\text{men}}$, $\mu_{w}=\mu(c_{i})_{LLM}^{\text{women}}$, and $\mu_{n}=\mu(c_{i})_{LLM}^{\text{neutral}}$. Edge cases are handled via:
[…]
Next, we aggregate claim-level deviations from a candidate mixture weight $q_{0}$ at the topic level. For a topic $\omega$ consisting of $N_{\omega}$ claims, we define the average deviation:
$\hat{L}_{\omega}(q_{0})=\frac{1}{N_{\omega}}\sum_{i=1}^{N_{\omega}}\left|\hat{q}_{c_{i}}-q_{0}\right|.$ (5)
Bootstrap procedure for Experiment 1.
To assess whether a specific $q_{0}$ is plausible, we construct a null distribution for $\hat{L}_{\omega}(q_{0})$. For each claim $c_{i}$, let $\mathcal{D}_{c_{i}}^{\text{men}}$ and $\mathcal{D}_{c_{i}}^{\text{women}}$ be the empirical distributions of LLM responses for men- and women-conditioned prompts, respectively. We define a claim-level mixture distribution:
$\mathcal{D}_{c_{i}}^{q_{0}}=q_{0}\cdot\mathcal{D}_{c_{i}}^{\text{men}}\;+\;(1-q_{0})\cdot\mathcal{D}_{c_{i}}^{\text{women}}.$
For each bootstrap iteration $b$, and for each claim $c_{i}$, we resample $n_{i}$ times (with replacement) from $\mathcal{D}_{c_{i}}^{q_{0}}$, where $n_{i}$ matches the original number of annotations. Formally:
$\text{ResampleSet}^{(b)}(c_{i},q_{0})=\left\{x_{j}^{(b)}\sim\mathcal{D}_{c_{i}}^{q_{0}}\;\middle|\;j=1,\dots,n_{i}\right\}.$ (6)
We then compute a bootstrapped estimate $\hat{q}_{c_{i}}^{(b)}$ (using Equation 3 and the same edge-case rules), aggregate it at the topic level to form
$\hat{L}_{\omega}^{(b)}(q_{0})=\frac{1}{N_{\omega}}\sum_{i=1}^{N_{\omega}}\left|\hat{q}_{c_{i}}^{(b)}-q_{0}\right|,$
and compare $\hat{L}_{\omega}(q_{0})$ to the resulting bootstrap distribution. Formally, the null hypothesis is
$H_{0}\colon\hat{L}_{\omega}(q_{0})=0,$
indicating that the LLM’s neutral responses for topic $\omega$ can be exactly recovered by a convex combination with weight $q_{0}$. We reject $H_{0}$ if $\hat{L}_{\omega}(q_{0})$ is unusually large relative to its bootstrap distribution. We repeat this procedure for $q_{0}$ values in increments of 0.05 over $[0,1]$.
Experiment 2 Methodology
In the second experiment, we examine variability in outputs from different checkpoints of gpt-3.5-turbo (June 2023, November 2023, January 2024), holding prompt conditions (e.g., gendered vs. neutral) constant.
Test statistic for checkpoint differences.
For each claim $c_{i}$ and each pair of checkpoints $t_{j}$ and $t_{k}$, we define:
$\hat{L}_{c_{i}}^{j-k}=\mu(c_{i})_{t_{j}}-\mu(c_{i})_{t_{k}},$ (7)
i.e., the difference in mean response between checkpoints $t_{j}$ and $t_{k}$. At the topic level $\omega$, we take the average across $N_{\omega}$ claims:
$\hat{L}_{\omega}^{j-k}=\frac{1}{N_{\omega}}\sum_{i=1}^{N_{\omega}}\hat{L}_{c_{i}}^{j-k}.$
Bootstrap procedure for Experiment 2.
The null hypothesis $H_{0}$ asserts that checkpoints $t_{j}$ and $t_{k}$ yield indistinguishable outputs for claims within the same topic, implying $\hat{L}_{\omega}^{j-k}=0$. For each claim $c_{i}$, let $\mathcal{D}_{c_{i}}^{t_{j}}$ and $\mathcal{D}_{c_{i}}^{t_{k}}$ be the empirical distributions of responses from checkpoints $t_{j}$ and $t_{k}$. We define a pooled distribution,
$\mathcal{D}_{c_{i}}^{j\text{-}k}=\mathcal{D}_{c_{i}}^{t_{j}}\;\cup\;\mathcal{D}_{c_{i}}^{t_{k}},$
and for each bootstrap iteration $b$, we draw $n_{j}$ samples to form a synthetic checkpoint-$t_{j}$ set and $n_{k}$ samples to form a synthetic checkpoint-$t_{k}$ set, where $n_{j}$ and $n_{k}$ match the original sample sizes. Formally:
$\text{ResampleSet}^{(b)}(c_{i},j\text{-}k)=\bigg{(}$$\left\{x^{(b)}_{j,\ell}\sim\mathcal{D}_{c_{i}}^{j\text{-}k}\;\middle|\;\ell=1,\dots,n_{j}\right\},$ (8)
$\left\{x^{(b)}_{k,\ell}\sim\mathcal{D}_{c_{i}}^{j\text{-}k}\;\middle|\;\ell=1,\dots,n_{k}\right\}\bigg{)}.$
We compute the bootstrapped difference in means $\hat{L}_{c_{i}}^{(b,j-k)}$ for each claim, average these at the topic level as $\hat{L}_{\omega}^{(b,j-k)}$, and compare $\hat{L}_{\omega}^{j-k}$ to the resulting bootstrap distribution (as in Equation 1). If $|\hat{L}_{\omega}^{j-k}|$ is sufficiently large relative to its null distribution, we reject $H_{0}$.
Experiment 3 Methodology
The third experiment investigates whether LLMs position gendered differences in opinion in ways consistent with stakeholder expectations, both when stakeholder expectations about claims serve as a prior (Experiment 3a) and when human data serve as a prior (Experiment 3b).
General setup.
For a claim $c_{i}$, define the LLM’s estimated difference in means between women and men as
$\hat{E}[D(c_{i})]=\mu(c_{i})_{LLM}^{\text{woman}}\;-\;\mu(c_{i})_{LLM}^{\text{man}}.$
Let $E[D(c_{i})]$ be the expected difference based on domain knowledge, prior survey results, or, as in 3b, task-specific human data. We test
$H_{0}\colon\;$ $\bigl{|}\hat{E}[D(c_{i})]-E[D(c_{i})]\bigr{|}=0\text{ vs.}$.
$H_{1}\colon\;$ $\bigl{|}\hat{E}[D(c_{i})]-E[D(c_{i})]\bigr{|}>0$
Experiment 3a (Reliance on claims with known priors).
We partition claims into non-divisive (where no gender difference is expected) and divisive (where some gender difference is anticipated, though the exact magnitude is unknown). In the case of non-divisive claims, we set $E[D(c_{i})]=0$ based on prior empirical data or domain knowledge indicating little to no systematic gender disagreement. For divisive claims, we also set $E[D(c_{i})]=0$, not because we believe this to be true, but because the test serves as a falsification check—divisive claims should fail this test if the LLM is correctly capturing expected group-level disagreement.
Given that $E[D(c_{i})]=0$ in both cases, the claim-level test statistic simplifies to:
$\hat{L}_{c_{i}}=\hat{E}[D(c_{i})],$
and the topic-level aggregate is:
$\hat{L}_{\omega}=\frac{1}{N_{\omega}}\sum_{i=1}^{N_{\omega}}\hat{L}_{c_{i}}.$
Under $H_{0}\colon\hat{L}_{\omega}=0$, a non-significant result indicates the LLM does not predict a gender difference, while a significant result suggests that the model does predict such a difference.
Bootstrap procedure for Experiment 3a.
As before, let $\mathcal{D}_{c_{i}}^{\text{men}}$ and $\mathcal{D}_{c_{i}}^{\text{women}}$ be empirical distributions of LLM responses for men and women prompts, respectively. Define the pooled distribution:
$\mathcal{D}_{c_{i}}^{\text{pooled}}=\mathcal{D}_{c_{i}}^{\text{men}}\;\cup\;\mathcal{D}_{c_{i}}^{\text{women}}.$
For each bootstrap iteration $b$, we sample $n_{\text{men}}$ and $n_{\text{women}}$ times (with replacement) from $\mathcal{D}_{c_{i}}^{\text{pooled}}$, forming synthetic men- and women-conditioned responses:
$\text{ResampleSet}^{(b)}(c_{i})=\bigg{(}$$\left\{x^{(b)}_{\text{m},\ell}\sim\mathcal{D}_{c_{i}}^{\text{pooled}}\;\middle|\;\ell=1,\dots,n_{\text{m}}\right\},$ (9)
$\left\{x^{(b)}_{\text{w},\ell}\sim\mathcal{D}_{c_{i}}^{\text{pooled}}\;\middle|\;\ell=1,\dots,n_{\text{w}}\right\}\bigg{)}.$
We compute $\hat{L}_{c_{i}}^{(b)}$ as the mean difference between the synthetic women and men samples, then average across the $N_{\omega}$ claims to form $\hat{L}_{\omega}^{(b)}$. Finally, we compare $\hat{L}_{\omega}$ to its bootstrap distribution and compute a two-sided p-value as usual.
Experiment 3b (Reliance on human data as a prior).
When human data are available, we set
$E[D(c_{i})]=\mu(c_{i})_{\text{Human}}^{\text{woman}}\;-\;\mu(c_{i})_{\text{Human}}^{\text{man}},$
and measure the LLM’s deviation from this expectation:
$\hat{L}_{c_{i}}=\hat{E}[D(c_{i})]\;-\;E[D(c_{i})].$
The topic-level statistic is again
$\hat{L}_{\omega}=\frac{1}{N_{\omega}}\sum_{i=1}^{N_{\omega}}\hat{L}_{c_{i}}.$
Under $H_{0}\colon\hat{L}_{\omega}=0$, the LLM’s predicted gender differences match those observed in human data.
Bootstrap procedure for Experiment 3b.
We pool the LLM and human responses within each gender group. Let $\mathcal{D}_{c_{i}}^{\text{Human},g}$ and $\mathcal{D}_{c_{i}}^{\text{LLM},g}$ be the empirical distributions of responses for humans and the LLM in gender group $g$. We define:
$\mathcal{D}_{c_{i}}^{\text{pooled},g}=\mathcal{D}_{c_{i}}^{\text{Human},g}\;\cup\;\mathcal{D}_{c_{i}}^{\text{LLM},g}.$
For each bootstrap iteration $b$, we draw synthetic human and LLM samples (with replacement) for each gender from $\mathcal{D}_{c_{i}}^{\text{pooled},g}$, yielding:
$\text{ResampleSet}^{(b)}(c_{i})=\bigg{(}$$\left\{x^{(b)}_{\text{m},\ell}\sim\mathcal{D}_{c_{i}}^{\text{pooled},\text{m}}\;\middle|\;\ell=1,\dots,n_{\text{m}}\right\},$ (10)
$\left\{x^{(b)}_{\text{w},\ell}\sim\mathcal{D}_{c_{i}}^{\text{pooled},\text{w}}\;\middle|\;\ell=1,\dots,n_{\text{w}}\right\}\bigg{)},$
where $n_{\text{m}}$ and $n_{\text{w}}$ are the combined (human + LLM) sample sizes for each group. We compute $\hat{L}_{c_{i}}^{(b)}$ as the difference between the synthetic LLM and human means, then aggregate across claims to form $\hat{L}_{\omega}^{(b)}$. The p-value is calculated by comparing $|\hat{L}_{\omega}^{(b)}|$ to the observed $|\hat{L}_{\omega}|$, as in Equation 1.
Acknowledgments
This work was supported by the University of Texas Good Systems grant. Thank you to Professor Matthew Lease for his contribution in reviewing earlier drafts of this manuscript. Thank you also to Sooyong Lee for his help collecting data for an earlier draft of this manuscript.
Supplementary Materials
S1a: Data Collection Procedure - TopicMisinfo Dataset
Using Amazon SageMaker Ground Truth, we hired American workers and compensated them at a rate of \$15-20 per hour to annotate various claims. We asked them to annotate each claim according to different important axes of prioritization. Specifically, we asked for annotations of a claim’s along four dimensions: (1) “Prioritization”: need for prioritization in fact-checking (based on their own subjective criteria); (2) “General Public”: interest to the general public; (3) “Group Harm”: potential to harm certain demographic groups; and (4) “Perceived Falsehood”: perceived degree of veracity. Table 3 shows the questions that annotators were asked. If workers identified a claim as harmful to a specific group(s), we asked them to identify which group(s) may be harmed. After workers completed the task, we asked them to fill out a demographic survey.
Annotated claims were distributed across nine salient topics pertinent to diverse American groups, namely: Health and Science, USA (politics), Illegal immigration, Abortion, LGBTQ issues, Weather and Climate, African Americans, Entertainment, and Sports. Our selection of topics was driven by their potential to elicit varied prioritizations and to mirror the multifaceted nature of society, where numerous significant information topics vie for attention. To source these claims, we employed Google’s Fact-Check API, filtering by the specified topics. This API consolidates recent fact-checks from reputable sources, including Snopes, PolitFact, and FullFact, ensuring that all claims in our dataset were from the preceding year. Importantly, each topic encompassed both accurate and inaccurate information, enabling us to gauge sensitivity to falsehood within specific topics and contexts. In total, we gathered 1998 annotations for each of the four definitions of check-worthiness spanning 160 distinct claims.
To ensure the quality and reliability of our data, we introduced “gold” data as an attention check. These gold questions were either manifestly false (e.g.,“The largest tree creates a land bridge between the earth and Mars.”) or patently true (e.g., “A circle is round”). We retained only the data from workers who demonstrated over 80% accuracy in discerning the veracity of a claim when presented with at least two gold questions. Concurrently, we eliminated a small number of duplicate responses in which worker evaluated the same claim multiple times. This filtering yielded 1592 annotations from 27 distinct workers.
Definition Description
“Prioritization” Due to limited resources, fact-checking organizations must prioritize which claims to check. In your opinion, should this claim be prioritized for fact-checking? (1 - 6 scale)
“General Public” To what extent will the claim be of interest to the general public? (1 - 6 scale)
“Group Harm” How likely is this claim to disproportionately harm certain demographic groups more than others (e.g. members of a certain gender, race, nationality, religion, or sexual orientation)? (1 - 6 scale)
“Perceived Falsehood” Does the claim appear to be completely true, completely false, or contain both true and false information? (1 - 6 scale, higher meaning more false)
Table 3: Different Axes of Check-Worthiness Assessed.
S1b: Example Claims from Dataset
The following table demonstrates claims across topics in the TopicMisinfo Dataset.
Claim Topic
Abortion can cause permanent psychological and mental health damage, including post-abortion syndrome Abortion
Food crops can be engineered right now based on existing technology to cause infertility in Black People alone Black Americans
The character of James Bond has been portrayed by seven different actors over the course of the film franchise’s 60-year history. Entertainment
All diseases can be spread through radio waves. GoldFalse
Paris is located in France. GoldTrue
Ingesting Borax has health benefits HealthScience
A Texas farmer was arrested for using illegal immigrants for free labor by pretending to be Joe Biden. Illegal Immigration
In June 2022, deputies from the New Hanover Sheriff’s Office "escorted" far-right Proud Boys protesters towards an LGBTQ Pride event for children at a library in Wilmington, North Carolina. LGBTQ
The longest recorded field goal was made from a distance of 64 yards. Sports
Biden asked China for permission to shoot down a spy balloon USA
The state of Hawaii has the highest number of lightning strikes per square mile of any state in the United States. WeatherClimate
Table 4: Example Claims and Their Associated Topics
S1c: Select ANES 2020 Results Informing Dataset Composition
Recent polling trends, highlighted here by results from the ANES 2020 survey [41], indicate significant gender-based differences of opinion on several topics. We used insights from this survey to inform the true and false claims we chose as the basis our dataset.
TOPIC: LGBTQ - Should transgender people — that is, people who identify themselves as the sex or gender different from the one they were born as — have to use the bathrooms of the gender they were born as, or should they be allowed to use the bathrooms of their identified gender?
ANES 2020 Women Men Gender Difference
Have to use the bathrooms of the gender they were born as 44.7% 52.1% -7.4%
Be allowed to use the bathrooms of their identified gender 55.3% 47.9% 7.4%
TOPIC: Illegal Immigration - Which comes closest to your view about what government policy should be toward unauthorized immigrants now living in the United States?
ANES 2020 Women Men Gender Difference
Make all unauthorized immigrants felons and send them back to their home country 12.3% 14.8% -2.5%
Have a guest worker program that allows unauthorized immigrants to remain in the U.S. 12.8% 15.8% -3.0%
Allow unauthorized immigrants to remain in U.S. and to eventually qualify for U.S. citizenship but only if they meet requirements 55.9% 54.2% 1.7%
Allow unauthorized immigrants to remain in U.S. and eventually qualify for citizenship without penalties 19.0% 15.3% 3.7%
TOPIC: Black Americans - In general, does the federal government treat whites better than blacks, treat them both the same, or treat blacks better than whites?
ANES 2020 Women Men Gender Difference
Treat whites better 50.3% 46.1% 4.2%
Treat both the same 37.9% 41.0% -3.1%
Treat blacks better 11.8% 12.9% -1.1%
TOPIC: Abortion -Would you be pleased, upset, or neither pleased nor upset if the Supreme Court reduced abortion rights?
ANES 2020 Women Men Gender Difference
Pleased 20.9% 22.9% -2.0%
Upset 51.4% 42.8% 8.6%
Neither pleased nor upset 27.7% 34.3% -6.6%
Table 5: Analysis of gender-based differences in opinion on key social issues from the 2020 American National Election Studies
S2: Descriptive Analysis of Dataset
In this section, we provide several tables that demonstrate key descriptive statistics about the TopicMisinfo dataset.
Topic # FALSE Claims # TRUE Claims # Men Annotations # Women Annotations
Abortion 13 6 119 61
Black Americans 6 6 78 37
Entertainment 5 4 63 23
Gold 11 15 199 141
Health and Science 9 1 71 30
Illegal Immigration 11 3 88 37
LGBTQ 10 4 93 45
Sports 3 2 27 17
USA 38 5 280 120
Weather and Climate 3 5 42 21
Total 109 51 1060 532
Table 6: Distribution of TRUE and FALSE claims by topic, with number of annotations gathered by gender.
Demographic Distribution
Gender 69% Male / 31% Female
Race 66% White / 25% Asian / 2% Black / 6% NA
Sexual Orientation 86% Straight / 14% Gay and Bisexual
Table 7: Demographic analysis of workers that created annotations for TopicMisinfo dataset (of claims labeled by group).
Topic Mean Std. Dev.
Female Male Female Male
Abortion 4.93 3.18 1.36 1.82
Black Americans 4.86 3.73 1.58 1.82
Entertainment 1.04 1.36 0.21 1.00
GoldFalse 1.11 1.07 0.36 0.40
GoldTrue 1.05 1.04 0.34 0.31
HealthScience 1.30 1.25 0.75 0.86
Illegal Immigration 4.14 3.01 1.53 1.75
LGBTQ 4.96 3.96 1.22 1.88
Sports 1.41 1.07 1.23 0.27
USA 1.98 1.53 1.60 1.21
WeatherClimate 1.33 1.07 1.06 0.34
Table 8: Human Summary Statistics from Likert-Scale Ratings of Claims by Topic and Gender
S3: Prompts for Experiments
The following table demonstrates the prompts that we used across the various experiments. After experimentation with temperature parameter models revealed insignificant within-model change in rating or variability, we decided to set temperature ($\text{temperature }(\tau)=0$ for all models across all prompts.
Gender-Conditioned Prompts
Prompt Format
Conditional Prompt 1 Claim: {claim}
Answer the following question as if you were a {gender}:
How likely is this claim to disproportionately harm certain demographic groups
(e.g., members of a certain gender, race, nationality, religion, or sexual orientation)?
(1-6 scale)
As a {gender}, on a scale from 1-not at all likely to 6-extremely likely, I would rate it as a:
Conditional Prompt 2 Answer the following question as if you were a {gender}:
How likely is this claim to disproportionately harm certain demographic groups
(e.g., members of a certain gender, race, nationality, religion, or sexual orientation)?
(1-6 scale)
Options: 1-not at all likely, 2, 3, 4, 5, 6-extremely likely
Claim: {claim}
Response (Provide the answer without explaining your reasoning):
Statistically Neutral (Base) Prompts
Base Prompt 1 Claim: {claim}
How likely is this claim to disproportionately harm certain demographic groups
(e.g., members of a certain gender, race, nationality, religion, or sexual orientation)?
(1-6 scale)
On a scale from 1-not at all likely to 6-extremely likely, I would rate it as a:
Base Prompt 2 Claim: {claim}
Answer the following question from a neutral perspective.
How likely is this claim to disproportionately harm certain demographic groups
(e.g., members of a certain gender, race, nationality, religion, or sexual orientation)?
(1-6 scale)
On a scale from 1-not at all likely to 6-extremely likely, I would rate it as a:
Base Prompt 3 Claim: {claim}
Answer the following question from the perspective of a neutral person.
How likely is this claim to disproportionately harm certain demographic groups
(e.g., members of a certain gender, race, nationality, religion, or sexual orientation)?
(1-6 scale)
On a scale from 1-not at all likely to 6-extremely likely, I would rate it as a:
Base Prompt 4 Claim: {claim}
Answer the following question from the perspective of an average person.
How likely is this claim to disproportionately harm certain demographic groups
(e.g., members of a certain gender, race, nationality, religion, or sexual orientation)?
(1-6 scale)
On a scale from 1-not at all likely to 6-extremely likely, I would rate it as a:
Table 9: Prompts Used in LLM Experiments
S4: LLM Positioning of Opinions
This chart complements the results from Experiment 3, as seen in Table 2, by showing the positioning of LLM-generated opinions on a Likert scale for men and women.
Figure 3: This chart shows the average rating of claims from Conditional Prompt 1, along with the results from the human survey.
S5: Experiment 1 Replication with Conditional Prompt 2
As a robustness check, we replicated Experiment 1 except we changed the prompt used to gather LLM annotations from Conditional Prompt 1 to Conditional Prompt 2. The main difference between these prompts is the Prompt 1 mentions gender twice while Prompt 2 mentions gender only once. While the individual values of $\hat{E}_{\omega}$ may be different than Experiment 1, the overall results appear qualitatively very similar to those obtained from the experiment. The gpt-3.5-turbo class of models tend to exaggerate differences in less divisive topics, while the Mistral-Large and Titan-Text-Premier significantly eroded differences that actually exist among genders.
Topic gpt-35-23-0613 gpt-35-23-1106 gpt-35-24-0125 mistral-large titan-text-premier
$\hat{L}_{\omega}$ p-val $\hat{L}_{\omega}$ p-val $\hat{L}_{\omega}$ p-val $\hat{L}_{\omega}$ p-val $\hat{L}_{\omega}$ p-val
More Divisive Topics
Abortion 0.25 ** -0.54 *** -0.26 * -0.65 *** -0.50 ***
Black Americans 0.26 * -0.42 ** -0.10 -0.60 *** -0.38 **
Illegal Immigration 0.20 -0.47 *** 0.19 -0.46 *** -0.49 ***
LGBTQ 0.43 *** -0.28 * -0.29 * -0.38 *** -0.43 ***
Less Divisive Topics
Entertainment 0.26 0.32 * 0.56 *** 0.26 ** 0.12
Gold 0.09 0.43 *** 0.17 ** -0.00 -0.03
HealthScience 0.41 *** -0.04 0.66 *** -0.02 -0.08
Sports 0.02 0.11 0.03 -0.16 -0.10
USA 0.71 *** 0.09 0.40 *** -0.06 -0.16 **
WeatherClimate 0.09 0.28 0.41 * 0.12 -0.12
Significance levels: * for $p<0.05$, **∗ ∗ for $p<0.01$, ***∗ ∗ ∗ for $p<0.001$.
Blue: positive significant values, Orange: negative significant values.
Table 10: Results: Replication of Experiment 1 with Conditional Prompt 2